# Magnitude homology of real hyperplane arrangements

### 04 Apr 2026 v1

@@ 2.4. Hyperplane arrangements
In this paper, all arrangements are real central hyperplane arrangements unless otherwise specified. Standard references are [27, 30]. Let $\mathcal{A}$ be an arrangement in $\mathbb{R}^{\ell}$. We denote by $L(\mathcal{A})=\{\cap_{H\in\mathcal{B}}H\mid\mathcal{B}\subseteq\mathcal{A}\}$ the intersection lattice of $\mathcal{A}$, ordered by reverse inclusions. An element of $L(\mathcal{A})$ is called a flat. Note that the maximal flat is $\hat{1}=\cap_{H\in\mathcal{A}}H=:\cap\mathcal{A}$, called the center of $\mathcal{A}$, and the minimal flat $\hat{0}$ is the ambient space $\mathbb{R}^{\ell}$. The arrangement is called essential if the center is the origin. The rank of a flat $X\in L(\mathcal{A})$ is defined as the codimension $\mathrm{rank}X=\mathrm{codim}X$. The rank of $\mathcal{A}$ is defined as $\mathrm{rank}\mathcal{A}=\mathrm{rank}(\cap\mathcal{A})$. We write $L_{k}(\mathcal{A})$ as the set of flats of rank $k$ and $L_{\leq k}(\mathcal{A})$ the set of flats of rank at most $k$. Using the Möbius function $\mu:L(\mathcal{A})\times L(\mathcal{A})\to\mathbb{Z}$, we define the characteristic polynomial of $\mathcal{A}$ by $\chi(\mathcal{A},t)=\sum_{X\in L(\mathcal{A})}\mu(\hat{0},X)t^{\dim X}$. It is classically known $\mathrm{Ch}(\mathcal{A})=|\chi(\mathcal{A},1)|$ [33].
The arrangement $\mathcal{A}$ determines a stratification of $\mathbb{R}^{\ell}$, the strata are called faces (covectors). Let $\mathcal{F}(\mathcal{A})$ be the poset of faces, ordered by closure inclusions. The unique minimal face is the center $\cap\mathcal{A}=\cap_{H\in\mathcal{A}}H$, and maximal faces are the chambers $\mathrm{Ch}(\mathcal{A})$. For faces $F,G\in\mathcal{F}(\mathcal{A})$, their meet $F\wedge G$ always exists, while their join $F\vee G$ may not exist. There is an order-reversing surjective map $s:\mathcal{F}(\mathcal{A})\to L(\mathcal{A})$ taking a face $F$ to its support $s(F)$, that is the minimal subspace of $\mathbb{R}^{\ell}$ containing $F$. We define the dimension of a face as the dimension of its support.
For $F\in\mathcal{F}(\mathcal{A})$, let $\mathcal{A}_{F}=\{H\in\mathcal{A}\mid F\subseteq H\}$. Also, for a flat $X\in L(\mathcal{A})$, let $\mathcal{A}_{X}=\{H\in\mathcal{A}\mid X\subseteq H\}$ and $\mathcal{A}^{X}=\{H\cap X\mid H\in\mathcal{A}\setminus\mathcal{A}_{X}\}$. Note that if $F,G\in\mathcal{F}(\mathcal{A})$ have the same support $X\in L(\mathcal{A})$, then $\mathcal{A}_{F}=\mathcal{A}_{G}=\mathcal{A}_{X}$.
@@
the number of faces whose support is $X$, which is also equal to $\mathrm{Ch}(\mathcal{A}^{X})$. If $X$ is not the ambient space $\hat{0}$$\mathbb{R}^{\ell}$, we also define the beta invariant of $X$ by
$\beta_{X}:=\beta(\mathcal{A}_{X}).$
Recall that the beta invariant of an arrangement is defined as
$\beta(\mathcal{A})=(-1)^{\mathrm{rank}\mathcal{A}-1}\chi^{\prime}(\mathcal{A},1)=|\chi^{\prime}(\mathcal{A},1)|.$
@@
The beta invariant $\beta_{X}$ counts the bounded chambers in the deconing of $\mathcal{A}_{X}$. In other words, let $F$ be a face whose support is $X$ and pick $H\in\mathcal{A}$ such that $F\subseteq H$, then $\beta_{X}=\frac{1}{2}\#\{C\in\mathrm{Ch}(\mathcal{A})\mid\overline{C}\cap H=\overline{F}\}$ (See [33, 2, 30]). Hence the product $c^{X}\beta_{X}$ is also a counting function
@@
For arrangement $\mathcal{A}$ in $V=\mathbb{R}^{\ell}$, we fix a defining form $\alpha_{H}\in V^{*}$ for each $H\in\mathcal{A}$ and consider the line segment $L_{H}=[-\alpha_{H},\alpha_{H}]$ in $V^{*}$. We write $Z(\mathcal{A})=\sum_{H\in\mathcal{A}}L_{H}$ as the Minkowski sum of the line segments and call it the zonotope of $\mathcal{A}$. The poset of nonempty faces of this polytope is anti-isomorphic to $\mathcal{F}(\mathcal{A})$. For a face $F\in\mathcal{F}(A)$, we say the corresponding face $z_{F}$ of $Z(\mathcal{A})$ dual to $F$. In particular, 0-faces (vertices) of $Z(\mathcal{A})$ are dual to chambers of $\mathcal{A}$.
With these $\alpha_{H}$ chosen, each face $F\in\mathcal{F}(\mathcal{A})$ can be represented by a sign vector $F=(F_{H})_{H\in\mathcal{A}}\subseteq\{+,-,0\}^{\mathcal{A}}$, where $F_{H}=\mathrm{sgn}(\alpha_{H}(F))$. Declare $0<\pm$ and this determines a partial order on sign vectors coordinate-wisely, which coincides with the face order of $\mathcal{F}(\mathcal{A})$. We define the Tits product on $\mathcal{F}(\mathcal{A})$ using sign vectors as follow. For $F,G\in\mathcal{F}(\mathcal{A})$, let $FG\in\mathcal{F}(A)$ be the sign vector
@@
This product does not depend on the choices of defining forms, since geometrically, $FG$ is the face we land in when we start from a relative interior point of $F$, move out a small enough distance towards a relative interior point of $G$ (see Figure 1).
@@
Figure 1. A geometric visualization of the non-commutative Tits product in the $A_{2}$ arrangement. Let $F$ be a ray (blue) and $G$ be a chamber (red). (a) The product FGFG represents taking a small step from $F$ towards $G$, which immediately lands in the adjacent chamber. (b) The product GFGF represents taking a small step from $G$ towards $F$. Because $G$ is an open chamber, a sufficiently small step remains entirely within $G$, so $GF=G$.
@@ 2.5. Tope graphs and graph-theoretic properties
For an arrangement $\mathcal{A}$ in $\mathbb{R}^{\ell}$, notice that the tope graph $\mathcal{T}(\mathcal{A})$ is isomorphic to the $1$-skeleton of $Z(\mathcal{A})$. For a face $F\in\mathcal{F}(A)$, we let $\mathcal{T}(\mathcal{A})_{F}$ be the induced subgraph of $\mathcal{T}(\mathcal{A})$ with vertices $\mathrm{Ch}(\mathcal{A})_{F}=\{C\in\mathrm{Ch}(\mathcal{A})\mid F\leq C\}$, or equivalently, $\mathcal{T}(\mathcal{A})_{F}$ is the $1$-skeleton of the face $z_{F}$ of $Z(\mathcal{A})$. We call $\mathcal{T}(\mathcal{A})_{F}$ the $F$-subgraph of $\mathcal{T}(\mathcal{A})$. We first prove a useful lemma.
@@
For $F\in\mathcal{F}(\mathcal{A})$, $\mathcal{T}(\mathcal{A})_{F}$ is a gated subgraph of $\mathcal{T}(\mathcal{A})$.
@@
We prove that the map $p_{F}:\mathcal{T}(\mathcal{A})\to\mathcal{T}(\mathcal{A})_{F}$ defined by $p_{F}(D)=FD$ is the gate projection, that is, we prove the equation
@@
holds for $D\in\mathrm{Ch}(\mathcal{A})$ and $C\in\mathrm{Ch}(\mathcal{A})_{F}$. In terms of sign vectors, the distance $d(X,Y)=\sum_{H\in\mathcal{A}}d(X,Y)_{H}$ is the Hamming distance (number of non-equal signs), where
@@
Let us prove both sides of equation (2.3) have the same evaluation on each $H\in\mathcal{A}$. If $H\notin\mathcal{A}_{F}$, then $F_{H}\neq 0$ and $(FD)_{H}=F_{H}$, also since $C\in\mathrm{Ch}(\mathcal{A})_{F}$, we have $C_{H}=F_{H}$. Therefore (2.3) evaluated at $H$ is $d(D,C)_{H}=d(D,FD)_{H}+0$, which is trivially true since both sides equal $1$ if $D_{H}\neq F_{H}$ and 0 if $D_{H}=F_{H}$. Next suppose $H\in\mathcal{A}_{F}$, we have $F_{H}=0$. Then $d(D,FD)_{H}=0$. The equation (2.3) evaluated at $H$ is $d(D,C)_{H}=0+d(FD,C)_{H}$, which is triviall true since $D_{H}=(FD)_{H}$. ∎
Here we record some important examples of tope graphs.
@@
Let $\mathcal{A}=Bl_{d}$ be the $d$-th Boolean arrangement, that is, the arrangement of $d$ coordinate hyperplanes of $\mathbb{R}^{d}$. It is clear that the tope graph $\mathcal{T}(\mathcal{A})$ is the whole hypercube graph $Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}\cong Q_{d}$.
@@
Let $r\geq 2$. An arrangement of $n$ hyperplanes in $\mathbb{R}^{r}$ is said to be generic if any subarrangement of $r$ hyperplanes has rank $r$. Boolean arrangements are generic. Generic arrangements are realizations of uniform matroids $U_{r,n}$. By abuse of notations, we use $U_{r,n}$ to denote a generic arrangement of $n$ hyperplanes in $\mathbb{R}^{r}$. For example, consider a generic arrangement $\mathcal{A}=U_{r,r+1}$. The tope graph $\mathcal{T}(\mathcal{A}_{W})$ is isomorphic to the induced subgraph of $Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}\cong Q_{r+1}$ with all vertices (sign vectors) except the all ++ vector and the all - vector.
@@
Consider a finite Coxeter system $(W,S)$ and the associated Coxeter arrangement $\mathcal{A}_{W}$. The zonotope $Z(\mathcal{A}_{W})$ is the $W$-permutohedron, and the tope graph $\mathcal{T}(\mathcal{A})$ is isomorphic to the Cayley graph of $(W,S)$.
Tope graphs have the following graph-theoretic properties. Recall that a chamber $C$ is simplicial if its intersection with the unit sphere of the ambient space is combinatorially a simplex. An arrangement is simplicial if all chambers all simplicial.
Proposition 2.30 (Propositions 4.2.15 and 4.4.8 of [4]).
Let $\mathcal{A}$ be an arrangement of $\mathrm{rank}\mathcal{A}=r$. The tope graph $\mathcal{T}=\mathcal{T}(\mathcal{A})$ has the following properties as an unlabeled graph.
@@
    For each vertex $C$, there is a unique vertex C-C such that $d(C,-C)=\mathrm{diam}(\mathcal{T})$.
@@
    The mapping $C\mapsto-C$ is a fixed point free automorphism of $\mathcal{T}$.
@@
    For each vertex $C$, $\deg(C)\geq r$, where the equality holds iff $C$ is simplicial.
@@
Recall that a graph is called $k$-regular if every vertex has degree $k$.
Corollary 2.31.
$\mathcal{A}$ is simplicial of rank $r$ if and only if $\mathcal{T}(\mathcal{A})$ is $r$-regular.
A matching of a graph is a collection of edges where no two edges share a common incident vertex. We say a matching is perfect if every vertex of the graph is incident to exactly one edge in the matching. The marriage theorem (see Theorem 2.1.2 of [9]) has a corollary (see Corollary 2.1.3 of [9]) asserting that every $k$-regular ($k\geq 1$) bipartite graph has a perfect matching.
Corollary 2.32.
If $\mathcal{A}$ is simplicial, then $\mathcal{T}(\mathcal{A})$ has a perfect matching.
Very recently, Hamiltonian cycles in tope graphs of simplicial and supersolvable arrangements are studied in [5, 20].

### 01 May 2026 v3

@@ 2.4. Hyperplane arrangements
The standard references are [28, 32]. Let $\mathcal{A}$ be an arrangement in $\mathbb{R}^{d}$. For two chambers $C,C^{\prime}\in\mathrm{Ch}(\mathcal{A})$, we write $S(C,C^{\prime})$ for the set of hyperplanes in $\mathcal{A}$ that separate $C$ and $C^{\prime}$. The tope graph $\mathcal{T}(\mathcal{A})$ of $\mathcal{A}$ is the graph with vertex set $\mathrm{Ch}(\mathcal{A})$ in which two vertices $C$ and $C^{\prime}$ are connected by an edge if $S(C,C^{\prime})$ consists of a single hyperplane. The graph metric $d$ on $\mathcal{T}(\mathcal{A})$ satisfies $d(C,C^{\prime})=\#S(C,C^{\prime})$. Note that $\mathcal{T}(\mathcal{A})$ is a partial cube, that is, it is isomorphic to an isometric subgraph of the hypercube graph $Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}$.
We denote by $L(\mathcal{A})=\{\cap_{H\in\mathcal{B}}H\mid\mathcal{B}\subseteq\mathcal{A}\}$ the intersection lattice of $\mathcal{A}$, ordered by reverse inclusion. An element of $L(\mathcal{A})$ is called a flat. Note that the maximal flat is $\hat{1}=\cap_{H\in\mathcal{A}}H=:\cap\mathcal{A}$, called the center of $\mathcal{A}$, and the minimal flat $\hat{0}$ is the ambient space $\mathbb{R}^{d}$. The arrangement is called essential if the center is the origin. The rank of a flat $X\in L(\mathcal{A})$ is defined as the codimension: $\operatorname{rank}X=\operatorname{codim}X$. The rank of $\mathcal{A}$ is defined as $\operatorname{rank}\mathcal{A}=\operatorname{rank}(\cap\mathcal{A})$. We write $L_{k}(\mathcal{A})$ for the set of flats of rank $k$ and $L_{\leq k}(\mathcal{A})$ for the set of flats of rank at most $k$. Using the Möbius function $\mu:L(\mathcal{A})\times L(\mathcal{A})\to\mathbb{Z}$, we define the characteristic polynomial of $\mathcal{A}$ by $\chi(\mathcal{A},t)=\sum_{X\in L(\mathcal{A})}\mu(\hat{0},X)t^{\dim X}$. It is classically known that $\mathrm{Ch}(\mathcal{A})=|\chi(\mathcal{A},-1)|$ [35].
The arrangement $\mathcal{A}$ determines a stratification of $\mathbb{R}^{d}$; the strata are called faces (covectors). Let $\mathcal{F}(\mathcal{A})$ be the poset of faces, ordered by closure inclusion. The unique minimal face is the center $\cap\mathcal{A}=\cap_{H\in\mathcal{A}}H$, and the maximal faces are the chambers $\mathrm{Ch}(\mathcal{A})$. For faces $F,G\in\mathcal{F}(\mathcal{A})$, their meet $F\wedge G$ always exists, while their join $F\vee G$ may not exist. There is an order-reversing surjective map $s:\mathcal{F}(\mathcal{A})\to L(\mathcal{A})$ taking a face $F$ to its support $s(F)$, that is, the minimal subspace of $\mathbb{R}^{d}$ containing $F$. We define the dimension of a face as the dimension of its support.
For a face $F\in\mathcal{F}(\mathcal{A})$, let $\mathcal{A}_{F}=\{H\in\mathcal{A}\mid F\subseteq H\}$. Also, for a flat $X\in L(\mathcal{A})$, let $\mathcal{A}_{X}=\{H\in\mathcal{A}\mid X\subseteq H\}$ and $\mathcal{A}^{X}=\{H\cap X\mid H\in\mathcal{A}\setminus\mathcal{A}_{X}\}$. Note that if $F,G\in\mathcal{F}(\mathcal{A})$ have the same support $X\in L(\mathcal{A})$, then $\mathcal{A}_{F}=\mathcal{A}_{G}=\mathcal{A}_{X}$.
@@
the number of faces whose support is $X$, which is also equal to $\mathrm{Ch}(\mathcal{A}^{X})$. If $X$ is not the ambient space $\hat{0}=\mathbb{R}^{d}$, we also define the beta invariant of $X$ by
$\beta_{X}:=\beta(\mathcal{A}_{X}),$
where the beta invariant of an arrangement is defined as
$\beta(\mathcal{A})=(-1)^{\operatorname{rank}\mathcal{A}-1}\chi^{\prime}(\mathcal{A},1)=|\chi^{\prime}(\mathcal{A},1)|.$
@@
The beta invariant $\beta_{X}$ counts the bounded chambers in the deconing of $\mathcal{A}_{X}$. In other words, let $F$ be a face whose support is $X$ and pick $H\in\mathcal{A}$ such that $F\subseteq H$. Then $\beta_{X}=\frac{1}{2}\#\{C\in\mathrm{Ch}(\mathcal{A})\mid\overline{C}\cap H=\overline{F}\}$ (see [35, 2, 32]). Hence the product $c^{X}\beta_{X}$ is also a counting function:
@@
For an arrangement $\mathcal{A}$ in $V=\mathbb{R}^{d}$, we fix a defining form $\alpha_{H}\in V^{*}$ for each $H\in\mathcal{A}$ and consider the line segment $L_{H}=[-\alpha_{H},\alpha_{H}]$ in $V^{*}$. We write $Z(\mathcal{A})=\sum_{H\in\mathcal{A}}L_{H}$ for the Minkowski sum of the line segments and call it the zonotope of $\mathcal{A}$. The poset of nonempty faces of this polytope is anti-isomorphic to $\mathcal{F}(\mathcal{A})$. For a face $F\in\mathcal{F}(\mathcal{A})$, we say that the corresponding face $z_{F}$ of $Z(\mathcal{A})$ is dual to $F$. In particular, the 0-faces (vertices) of $Z(\mathcal{A})$ are dual to the chambers of $\mathcal{A}$.
With these $\alpha_{H}$ chosen, each face $F\in\mathcal{F}(\mathcal{A})$ can be represented by a sign vector $F=(F_{H})_{H\in\mathcal{A}}\in\{+,-,0\}^{\mathcal{A}}$, where $F_{H}=\mathrm{sgn}(\alpha_{H}(F))$. Declaring $0<\pm$ determines a partial order on the set of sign vectors, which coincides with the face order of $\mathcal{F}(\mathcal{A})$. We define the Tits product on $\mathcal{F}(\mathcal{A})$ using sign vectors as follows. For $F,G\in\mathcal{F}(\mathcal{A})$, let $FG\in\mathcal{F}(\mathcal{A})$ be the sign vector
@@
This product does not depend on the choice of defining forms: geometrically, $FG$ is the face reached by starting from a relative interior point of $F$ and moving a sufficiently small distance toward a relative interior point of $G$ (see Figure 1).
@@
Figure 1. A geometric visualization of the non-commutative Tits product in the $A_{2}$ arrangement. Let $F$ be a ray (blue) and $G$ be a chamber (red). (a) The product FGFG represents taking a small step from $F$ toward $G$, which immediately lands in the adjacent chamber. (b) The product GFGF represents taking a small step from $G$ toward $F$. Because $G$ is an open chamber, a sufficiently small step remains entirely within $G$, so $GF=G$.
@@ 2.5. Tope graphs and graph-theoretic properties
For an arrangement $\mathcal{A}$ in $\mathbb{R}^{d}$, note that the tope graph $\mathcal{T}(\mathcal{A})$ is isomorphic to the $1$-skeleton of $Z(\mathcal{A})$. For a face $F\in\mathcal{F}(\mathcal{A})$, we let $\mathcal{T}(\mathcal{A})_{F}$ be the induced subgraph of $\mathcal{T}(\mathcal{A})$ with vertices $\mathrm{Ch}(\mathcal{A})_{F}=\{C\in\mathrm{Ch}(\mathcal{A})\mid F\leq C\}$; equivalently, $\mathcal{T}(\mathcal{A})_{F}$ is the $1$-skeleton of the face $z_{F}$ of $Z(\mathcal{A})$. We call $\mathcal{T}(\mathcal{A})_{F}$ the $F$-subgraph of $\mathcal{T}(\mathcal{A})$. We first prove a useful lemma.
@@
For $F\in\mathcal{F}(\mathcal{A})$, the subgraph $\mathcal{T}(\mathcal{A})_{F}$ of $\mathcal{T}(\mathcal{A})$ is gated with gate projection $D\mapsto FD$.
@@
We must show that the equality
@@
holds for $D\in\mathrm{Ch}(\mathcal{A})$ and $C\in\mathrm{Ch}(\mathcal{A})_{F}$. In terms of sign vectors, the distance $d(X,Y)=\sum_{H\in\mathcal{A}}d(X,Y)_{H}$ is the Hamming distance (the number of unequal signs), where
@@
We prove that both sides of equation (2.4) have the same contribution on each $H\in\mathcal{A}$. If $H\notin\mathcal{A}_{F}$, then $F_{H}\neq 0$ and $(FD)_{H}=F_{H}$; also, since $C\in\mathrm{Ch}(\mathcal{A})_{F}$, we have $C_{H}=F_{H}$. Therefore (2.4) evaluated at $H$ is $d(D,C)_{H}=d(D,FD)_{H}+0$, which is true since both sides equal $1$ if $D_{H}\neq F_{H}$ and 0 if $D_{H}=F_{H}$. Next suppose $H\in\mathcal{A}_{F}$. Then $F_{H}=0$ and $d(D,FD)_{H}=0$. The equation (2.4) evaluated at $H$ is $d(D,C)_{H}=0+d(FD,C)_{H}$, which is true since $D_{H}=(FD)_{H}$. ∎
We record some important examples of tope graphs.
@@
Let $\mathcal{A}=Bl_{d}$ be the $d$-th Boolean arrangement, that is, the arrangement of $d$ coordinate hyperplanes in $\mathbb{R}^{d}$. Then the tope graph $\mathcal{T}(\mathcal{A})$ is the whole hypercube graph $Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}\cong Q_{d}$.
@@
Let $r\geq 2$. An arrangement of $n$ hyperplanes in $\mathbb{R}^{r}$ is said to be generic if any subarrangement of $r$ hyperplanes has rank $r$. Boolean arrangements are generic. Generic arrangements are realizations of uniform matroids $U_{r,n}$. By abuse of notation, we use $U_{r,n}$ to denote a generic arrangement of $n$ hyperplanes in $\mathbb{R}^{r}$. For example, consider a generic arrangement $\mathcal{A}=U_{r,r+1}$. The tope graph $\mathcal{T}(\mathcal{A})$ is isomorphic to the induced subgraph of $Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}\cong Q_{r+1}$ obtained by removing the all-++ vector and the all-- vector.
@@
Consider a finite Coxeter system $(W,S)$ and the associated Coxeter arrangement $\mathcal{A}_{W}$. The zonotope $Z(\mathcal{A}_{W})$ is the $W$-permutahedron, and the tope graph $\mathcal{T}(\mathcal{A}_{W})$ is isomorphic to the Cayley graph of $(W,S)$.
Tope graphs have the following graph-theoretic properties. Recall that a chamber $C$ is simplicial if it is a simplicial cone. An arrangement is simplicial if all its chambers are simplicial.
Proposition 2.33 ([4, Propositions 4.2.15 and 4.4.8]).
Let $\mathcal{A}$ be an arrangement of $\operatorname{rank}\mathcal{A}=r$. The tope graph $\mathcal{T}=\mathcal{T}(\mathcal{A})$ has the following properties as an unlabeled graph.
@@
    For each vertex $C$, there is a unique vertex C-C such that $d(C,-C)=\operatorname{diam}(\mathcal{T})$.
@@
    The mapping $C\mapsto-C$ is a fixed-point-free automorphism of $\mathcal{T}$.
@@
    For each vertex $C$, we have $\deg(C)\geq r$, where equality holds if and only if $C$ is simplicial.